A random sample has a mean of 16.5 and variance of 16, with n = 35 . The null hypothesis Ho : ? = 20 is to be tested against the alternative hypothesis Ha : ? ? 20 with 99% confidence level. What is the value of the test statistics ? Write your answer with 2 decimal places.
Answer #1

Solution :

Given that,

\bar x = 16.5

s2 = 16

s = 4

\mu = 38 and n = 35

t = (\bar x - \mu) / s / \sqrtn = (16.5 - 20) / 4 / \sqrt35 = -5.18

Test statistic = -5.18